# Advancing the Understanding and Evaluation of AR-Generated Scenes: When Vision-Language Models Shine and Stumble

### v2

@@ 4.3 User Study
To compare human performance with VLMs in identifying and describing AR scenarios, we conducted a user study approved by the Duke University Campus Institutional Review Board (protocol number: 2020-0292). We recruited five participants from the Duke community, comprising four males and one female, aged between 23 and 29 years. Participants were informed that they would be shown a mix of AR and non-AR images and were tasked with providing responses based on the prompt T. To facilitate the study, we developed a user interface in Jupyter Notebook that displayed images at a size of 8 × 8 inches (20.32 × 20.32 cm), adjusted to fit the height of a laptop screen, allowing the entire image to be viewed clearly without scrolling. Researchers recorded participants’ responses through the interface by clicking buttons to label their evaluation outcomes.
@@
Description Performance: The average true positive rate for description ($TPR_{D}$) across medium and hard levels exceeded the best $TPR_{D}$ of VLMs by 7.7% and 27.1%, respectively, while falling only 1.1% short at the easy level. The superior performance at medium and hard levels can likely be attributed to humans’ ability to learn from observed images and apply reasoning skills. However, despite their enhanced reasoning, humans were prone to fatigue after reviewing approximately 100 images. Furthermore, we observed that the overall decline in human performance with increasing complexity levels closely mirrored the patterns exhibited by VLMs.
Factors Influencing Performance: Recognition of AR content among participants was influenced by various factors, such as object texture, size, lighting, shadows, and positioning. Across the study, participants misclassified virtual objects as real or vice versa in 7.6% to 13.8% of the images, amounting to 24 to 44 misclassifications per participant. Interestingly, nearly all of these misclassified images overlapped with those where VLMs also struggled, indicating similarities in the challenges faced by humans and VLMs. Moreover, 2 participants showed a tendency to focus on objects that were closer and centrally positioned in the image, particularly during the later stages of the study. This finding provides guidance for fine-tuning VLMs to prioritize closer and central areas of the image, aligning their attention more closely with human perception.
@@ 5 Conclusion and Future Work
In this work, we evaluate VLMs for AR scene understanding using the DiverseAR dataset, the first AR dataset to assess VLMs’ ability to identify and describe virtual content. Our results reveal that while VLMs perform well in detecting prominent AR elements, they struggle with seamlessly integrated virtual content, especially under realistic lighting, shadows, and physical plausibility conditions.
Building on this study, we are enhancing the DiverseAR dataset by including a larger collection of AR images, each annotated with quantifiable properties such as virtual object size, contrast, and position, while also collecting more human perception data for comparison. To further refine VLM performance, we are exploring AR-specific fine-tuning methods to align model capabilities with human perception, focusing on factors like shadow realism, lighting consistency, and object alignment.
Beyond image-level assessment, we aim to expand our evaluation to include quality factors related to temporal data modalities, such as the drift and jitter of virtual content. Building on the success of our previous non-VLM-based approaches [10, 19], we will explore VLM-based methods to effectively analyze and address these challenges.

### v3

@@ 4.3 User Study
To compare human performance with VLMs in identifying and describing AR scenarios, we conducted a user study approved by the Duke University Campus Institutional Review Board (protocol number: 2020-0292). We recruited five participants from the Duke community, comprising four males and one female, aged between 23 and 29 years. Participants were informed that they would be shown a mix of AR and non-AR images and were tasked with providing responses based on the prompt T. To facilitate the study, we developed a custom Jupyter Notebook interface that displayed each image at 8 × 8 inches (20.32 × 20.32 cm), adjusted to fit the height of a laptop screen, allowing the entire image to be viewed clearly without scrolling. Participants’ responses were recorded through the interface by clicking buttons to label their evaluation outcomes.
@@
Description Performance: The average True Positive Rate for Description ($TPR_{D}$) across medium and hard levels exceeded the best $TPR_{D}$ of VLMs by 7.7% and 27.1%, respectively, while falling only 1.1% short at the easy level. The superior performance at medium and hard levels can likely be attributed to humans’ ability to learn from observed images and apply reasoning skills. Despite their enhanced reasoning, humans were prone to fatigue after reviewing approximately 100 images. Furthermore, we observed that the overall decline in human performance with increasing complexity levels closely mirrored the pattern exhibited by VLMs.
Factors Influencing Performance: Recognition of AR content among participants was influenced by various factors, such as object texture, size, lighting, shadows, and positioning. Across the study, participants misclassified virtual objects as real or vice versa in 7.6% to 13.8% of the images, amounting to 24 to 44 misclassifications per participant. Interestingly, most misclassified images overlapped with those where VLMs also struggled, indicating similarities in the challenges faced by humans and VLMs. Moreover, 2 participants showed a tendency to focus on objects that were closer and centrally positioned in the image, particularly during the later stages of the study. This finding provides guidance for fine-tuning VLMs to prioritize closer and central areas of images, aligning their attention more closely with human perception.
@@ 5 Conclusion and Future Work
In this work, we evaluate VLMs for AR scene understanding using the DiverseAR dataset, the first AR dataset to assess VLMs’ ability to identify and describe virtual content. Our results reveal that while VLMs perform well in detecting prominent AR elements, they struggle with seamlessly integrated virtual content, especially under conditions of realistic lighting, shadows, and physical plausibility.
Building on this study, we are enhancing the DiverseAR dataset by gathering a larger collection of AR images, each annotated with quantifiable properties such as virtual object size, contrast, and position, while also collecting human perception data for comparison. To further refine VLM performance, we are exploring AR-specific fine-tuning methods to align model capabilities with human perception, focusing on factors like shadow realism, lighting consistency, and object alignment.
Beyond image-level assessment, we aim to expand our evaluation to include quality factors associated with temporal data modalities, such as the drift and jitter of virtual content. Building on the success of our previous non-VLM-based approaches [10, 19], we will explore VLM-based methods to further analyze and assess these factors.